### v2

@@ 1 Introduction
With a piecewise linear representation, we can embed the neural network in the optimization model using a Mixed-Integer Linear Programming (MILP) formulation. However, such formulations range from being small but having a weak linear relaxation to having a stronger relaxation but being prohibitivelly large [61], which makes it difficult to use an off-the-self MILP solver. That motivated studies on how to tackle such models, ranging from reformulation [61, 107, 123, 76] to methods for obtaining smaller big M coefficients through activation bounds [19, 33, 41, 74, 123, 6, 143, 57, 114] and parameter rescaling [98], activation inferences for fixing variables and limiting search space [120, 138, 11, 108], cutting planes [2], and heuristic solutions [96, 121]. Another—perhaps overlooked—approach is using neural networks that are smaller [15] and sparser [106, 138].
@@ 4.2 Results and Analysis
    The criteria of what connections to prune (such as MP vs. RP) has a significant effect on the results (xii), and may help extending this approach to neural networks with larger inputs and more layers (xiii):
@@
      The benefit of our approach is more significant in neural networks with smaller input size or smaller depth. On the other hand, layer width does not affect results in a clearly monotonic way.
@@ 5.2 Results and Analysis
Our approach based on Algorithm 2 generally produced better results for the instances that would be typically regarded as harder to solve for having a larger number of linear regions [121]. The results are not as favorable as they were for network verification, for which reason we use less space reporting and analyzing them. Nevertheless, they have some interesting similarities and differences when compared to those in Section 4. Table 3 summarizes the number of networks for which we found better values indirectly with Algorithm 2 according to input size, number of layers (network depth), and neurons per layer (layer width).

### v3

@@ 1 Introduction
With a piecewise linear representation, we can embed the neural network in the optimization model using a Mixed-Integer Linear Programming (MILP) formulation. However, such formulations range from being small but having a weak linear relaxation to having a stronger relaxation but being prohibitivelly large [63], which makes it difficult to use an off-the-self MILP solver. That motivated studies on how to tackle such models, ranging from reformulation [63, 107, 124, 76] to methods for obtaining smaller big M coefficients through activation bounds [19, 33, 41, 74, 124, 6, 144, 57, 114] and parameter rescaling [98], activation inferences for fixing variables and limiting search space [120, 139, 11, 108], cutting planes [2], and heuristic solutions [96, 121, 122]. Another—perhaps overlooked—approach is using neural networks that are smaller [15] and sparser [106, 139].
@@ 3.2.2 Function Maximization
Algorithm 2 Heuristic for maximizing the output of a (dense) neural network $\mathcal{D}$ by trying to solve the same problem on its sparse counterpart $\mathcal{S}$
@@
In both cases, we assume that the algorithms will be run up to a reasonable time limit, so that we would not be done solving the sparse model but would have obtained and evaluated a sufficient number of solutions in the dense model.
@@ 4.2 Results and Analysis
    Pruning criteria (such as MP vs. RP) affect the results significantly (xii), and may help tackling networks with larger inputs and more layers (xiii):
@@
      The benefit is more significant in neural networks with smaller input size or smaller depth, but less affected by layer width.
@@ 5.2 Results and Analysis
Our approach based on Algorithm 2 generally produced better results for the instances that would be typically regarded as harder to solve for having a larger number of linear regions [121]. These are modest results in comparison with Section 4, but with noteworthy similarities and differences. Table 3 summarizes the number of networks for which we found better values with Algorithm 2 by input size, number of layers (network depth), and neurons per layer (layer width).